-- RunaBriguglio - 05 Jul 2011

Coils test by Mario

found not working coils on acts 645 mir=454dsp, 582 mir=561dsp and 603 mir. 454 dsp that was replaced two days ago, 603 mirr never gave problems. after some verifications, we understood that 454 is ok but its DSP is not, and the same for 603mirr.

act replacement

act SN replaced with
561 1251 705
454   original

dsp replacement

after

crate dsp SN replaced with
1 4 247 143
4 0 287 288

Mario (MG) left the mountain.

Michael Wagner (LBTO) arrive on the mountain.

Sensors mapping

bcucoolerin1temp (bcu1, crate4) is 77°, error!

Problem with the pump of cooling: no flux. The work stops, Jeff will manage that with Joy tomorrow morning.
Topic revision: r6 - 06 Jul 2011, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback